# arXiv:2602.23696v1[cs.LG] 27 Feb 2026

### v1

@@ Abstract
We analyze the cumulative parameter trajectory of transformer training under AdamW and identify a stable, low-dimensional drift direction that captures the majority of long-horizon displacement from initialization. This “backbone” accounts for 60–80% of cumulative drift across blocks and remains temporally stable throughout training. Strikingly, instantaneous per-batch gradients are nearly isotropic relative to this direction, and it is orthogonal to top Fisher curvature modes. Instead, the backbone emerges from optimizer integration: momentum amplifies weak but temporally coherent gradient bias, while adaptive normalization suppresses incoherent transverse fluctuations.
Replacing AdamW with SGD-family optimizers eliminates this structure. Even at matched validation loss, SGD trajectories remain nearly colinear and fail to develop the multi-dimensional slow–fast geometry observed under AdamW, demonstrating that the backbone is optimizer-induced rather than loss-landscape-determined.
Oscillatory regime switching between competing objectives occurs primarily in directions transverse to the backbone. Reheating experiments show that transverse modes can be transiently re-excited from late-training checkpoints without erasing accumulated backbone drift, consistent with a slow–fast decomposition of training dynamics.
These results shift attention from instantaneous gradient geometry to cumulative trajectory structure and provide a concrete empirical characterization of optimizer-induced implicit bias in transformer training.
@@ 1 Introduction
In this work, we study the global parameter trajectory of transformer training under AdamW (Loshchilov and Hutter, 2019). Rather than analyzing local curvature or single-step gradients, we examine cumulative displacement from initialization across checkpoints. We find that training admits a stable low-dimensional drift direction—which we term the backbone—that captures the majority of cumulative parameter movement. Across blocks and seeds, the first principal component of uncentered trajectory PCA explains 60–80% of total drift, and this direction remains nearly fixed throughout training.
Crucially, the backbone is not aligned with instantaneous gradient directions and is nearly orthogonal to leading Fisher curvature modes (Martens, 2020). Per-batch gradients are close to isotropic relative to the backbone. However, the optimizer-integrated update—after momentum accumulation and adaptive per-parameter normalization (Kingma and Ba, 2015)—exhibits strong alignment with it. This demonstrates that the backbone is not a property of the loss landscape alone, but an emergent property of optimizer dynamics.
@@ 3.2 Temporal Stability of the Backbone
The mean of $c(w)$ across all windows is 0.9970.9970.9980.998 for both seeds, indicating that the backbone direction is essentially fixed from early training onward. It is not an artifact of averaging over distinct dynamical phases.
@@ 3.3 Backbone–Residual Decomposition
4 Mechanism: How the Backbone Emerges
This section presents the central result. We show that the backbone is an emergent property of optimizer integration, not of instantaneous gradient alignment.
4.1 The Puzzle: Per-Batch Gradients Do Not Align
An intuitive expectation is that the backbone arises because gradients consistently point along it. This is false.
At each checkpoint, we compute the per-batch combined-loss gradient $\mathbf{g}_{t}=\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta}(t))$ and measure its cosine similarity with $\mathbf{v}_{\mathrm{b}}$. Across 16 mini-batches per checkpoint:
$\lvert\cos\angle(\mathbf{g}_{t},\,\mathbf{v}_{\mathrm{b}})\rvert\;\approx\;0.008\text{--}0.012.$ (10)
To understand why, recall that for two independent random unit vectors $\mathbf{a},\mathbf{b}\in\mathbb{R}^{D}$, the expected absolute cosine similarity is
$\operatorname{\mathbb{E}}\bigl[\lvert\langle\mathbf{a},\,\mathbf{b}\rangle\rvert\bigr]\;=\;\sqrt{\frac{2}{\pi\,D}}.$ (11)
For $D=3.1\times 10^{6}$, this yields $\approx 4.5\times 10^{-4}$. The observed $\lvert\cos\rvert\approx 0.01$ is modestly above this (due to gradient structure), but still at the noise floor.
Proposition 1 (Noise-floor scaling).
For a gradient vector $\mathbf{g}\in\mathbb{R}^{D}$ with effective dimensionality $d_{\mathrm{eff}}\ll D$ (i.e., the gradient energy concentrates on $d_{\mathrm{eff}}$ directions), the expected projection onto a fixed unit vector $\mathbf{v}_{\mathrm{b}}$ scales as
$\operatorname{\mathbb{E}}\bigl[\lvert\langle\hat{\mathbf{g}},\,\mathbf{v}_{\mathrm{b}}\rangle\rvert\bigr]\;\sim\;\frac{1}{\sqrt{d_{\mathrm{eff}}}},$ (12)
where $\hat{\mathbf{g}}=\mathbf{g}/\lVert\mathbf{g}\rVert$. Since gradients in deep networks typically have $d_{\mathrm{eff}}\gg 1$, the alignment with any fixed direction is small.
No individual gradient step “points along” the backbone.
4.2 Optimizer Integration Resolves It
The backbone is not produced by strong instantaneous gradient alignment. Per-batch gradients are nearly isotropic with respect to the backbone direction, and their cosine alignment is only marginally above random. Instead, the backbone emerges from the interaction between optimizer dynamics and temporal coherence.
Under AdamW (Loshchilov and Hutter, 2019), the applied parameter update is not the raw gradient but a momentum-accumulated, variance-normalized step with weight decay. At step $t$, the first and second moment estimates are
$\mathbf{m}_{t}$ $\;=\;\beta_{1}\,\mathbf{m}_{t-1}\;+\;(1-\beta_{1})\,\mathbf{g}_{t},$ (13)
$\mathbf{v}_{t}$ $\;=\;\beta_{2}\,\mathbf{v}_{t-1}\;+\;(1-\beta_{2})\,\mathbf{g}_{t}^{2},$ (14)
where $\mathbf{g}_{t}^{2}$ denotes element-wise squaring, $\beta_{1}=0.9$, and $\beta_{2}=0.95$. With bias correction ($\hat{\mathbf{m}}_{t}$, $\hat{\mathbf{v}}_{t}$), the effective parameter update (including decoupled weight decay with coefficient $\mu$) is
$\boxed{\bm{\theta}_{t+1}\;=\;\bm{\theta}_{t}\;-\;\eta_{t}\left(\frac{\hat{\mathbf{m}}_{t}}{\sqrt{\hat{\mathbf{v}}_{t}}+\epsilon}\;+\;\mu\,\bm{\theta}_{t}\right),}$ (15)
where $\eta_{t}$ is the learning rate at step $t$ and $\epsilon=10^{-8}$. The net update direction is therefore
$\mathbf{u}_{t}\;=\;-\frac{\hat{\mathbf{m}}_{t}}{\sqrt{\hat{\mathbf{v}}_{t}}+\epsilon}\;-\;\mu\,\bm{\theta}_{t}.$ (16)
Three properties of this update conspire to create the backbone:
(i) Momentum integration of signed bias.
The exponential moving average in eq.˜13 acts as a low-pass filter with time constant $\tau_{m}=-1/\ln\beta_{1}\approx 9.5$ steps. Even when the instantaneous projection $\langle\mathbf{g}_{t},\,\mathbf{v}_{\mathrm{b}}\rangle$ is small, if its sign exhibits a weak but persistent bias across steps, the momentum term $\mathbf{m}_{t}$ integrates this bias. Concretely, suppose the $i$-th coordinate of the gradient has a mean $\mu_{i}$ and variance $\sigma_{i}^{2}$ across time. Then
$\operatorname{\mathbb{E}}[(\mathbf{m}_{t})_{i}]\;\approx\;\mu_{i},\qquad\operatorname{Var}[(\mathbf{m}_{t})_{i}]\;\approx\;\frac{1-\beta_{1}}{1+\beta_{1}}\,\sigma_{i}^{2},$ (17)
so the signal-to-noise ratio of the momentum is
$\mathrm{SNR}_{i}\;=\;\frac{\lvert\mu_{i}\rvert}{\sqrt{\operatorname{Var}[(\mathbf{m}_{t})_{i}]}}\;\approx\;\frac{\lvert\mu_{i}\rvert}{\sigma_{i}}\cdot\sqrt{\frac{1+\beta_{1}}{1-\beta_{1}}}\;=\;\frac{\lvert\mu_{i}\rvert}{\sigma_{i}}\cdot\sqrt{19},$ (18)
amplifying the per-coordinate SNR by a factor of $\sqrt{19}\approx 4.4$. Coordinates aligned with the backbone (where $\mu_{i}$ is nonzero due to the persistent LM gradient bias) benefit from this amplification.
(ii) Adaptive normalization suppresses transverse variance.
The denominator $\sqrt{\hat{\mathbf{v}}_{t}}+\epsilon$ in eq.˜15 normalizes each coordinate by its root-mean-square gradient magnitude. Coordinates with large but incoherent gradients (high $\sigma_{i}$, low $\mu_{i}$—typical of transverse directions) are divided by a large value, suppressing their contribution to the update. Coordinates with smaller but coherent gradients (moderate $\sigma_{i}$, nonzero $\mu_{i}$—typical of backbone-aligned directions) are normalized by a smaller value, preserving their contribution. Formally, the effective update along coordinate $i$ scales as
$\frac{(\hat{\mathbf{m}}_{t})_{i}}{\sqrt{(\hat{\mathbf{v}}_{t})_{i}}+\epsilon}\;\approx\;\frac{\mu_{i}}{\sqrt{\mu_{i}^{2}+\sigma_{i}^{2}}}\;=\;\frac{\mathrm{SNR}_{i}^{0}}{\sqrt{1+(\mathrm{SNR}_{i}^{0})^{2}}},$ (19)
where $\mathrm{SNR}_{i}^{0}=\mu_{i}/\sigma_{i}$. This is a sign-preserving squashing function: it amplifies directions with high SNR (backbone) relative to those with low SNR (transverse).
(iii) Cumulative coherence vs. instantaneous isotropy.
In $D=3.1\times 10^{6}$ dimensions, each gradient step has enormous freedom. Large transverse components can dominate the norm of each update while still cancelling over time if their directions fluctuate. A much smaller but temporally coherent component can therefore dominate cumulative displacement. If the per-step transverse displacement has variance $\sigma_{\perp}^{2}$ per dimension and the backbone displacement has magnitude $\delta_{\parallel}$ per step, then after $T$ steps the expected squared displacements are
$\lVert\Delta_{\parallel}\rVert^{2}\;=\;(T\,\delta_{\parallel})^{2}\;=\;T^{2}\,\delta_{\parallel}^{2},\qquad\operatorname{\mathbb{E}}[\lVert\Delta_{\perp}\rVert^{2}]\;=\;T\,(D-1)\,\sigma_{\perp}^{2},$ (20)
where the backbone displacement grows linearly ($\propto T$) while the transverse displacement grows as a random walk ($\propto\sqrt{T}$). The backbone fraction therefore approaches 1 as $T\to\infty$:
$f_{\mathrm{b}}\;\approx\;\frac{T^{2}\,\delta_{\parallel}^{2}}{T^{2}\,\delta_{\parallel}^{2}+T\,(D-1)\,\sigma_{\perp}^{2}}\;\xrightarrow{T\to\infty}\;1.$ (21)
This explains why $\rho_{1}\approx 0.80$ with only $T=51$ checkpoints: even a weak backbone bias, amplified by momentum and adaptive normalization, is sufficient to dominate cumulative drift.
4.3 Evidence: Update-Direction Alignment
We compute the 200-step update direction
$\mathbf{u}(t)\;=\;\bm{\theta}(t)-\bm{\theta}(t-200),$ (22)
which reflects the net effect of all optimizer operations (momentum, adaptive learning rates, weight decay, gradient clipping) accumulated over 200 steps. We measure the cosine similarity with the backbone:
$C(t)\;=\;\cos\angle\bigl(\mathbf{u}(t),\,\mathbf{v}_{\mathrm{b}}\bigr)\;=\;\frac{\langle\mathbf{u}(t),\,\mathbf{v}_{\mathrm{b}}\rangle}{\lVert\mathbf{u}(t)\rVert}.$ (23)
@@ 4.3 Evidence: Update-Direction Alignment
Phase Block 0 Blocks 1–7 (median) Interpretation
Early ($t<2000$) 0.27 (peak 0.34) 0.20–0.21 Strong alignment
Mid ($2000\leq t\leq 6000$) 0.054 0.054–0.076 Weakened at $\lambda$-transition
Late ($t>6000$) 0.110 0.061–0.143 Moderate, sign reversed
@@
Figure 2: The linchpin: update vs. gradient alignment with the backbone (seed 42, Block 0). Per-batch gradient alignment $\lvert\cos(\mathbf{g}_{t},\mathbf{v}_{\mathrm{b}})\rvert$ (blue circles with error bars) hovers at the noise floor (${\sim}0.005$), while the 200-step update alignment $\lvert\cos(\mathbf{u}_{t},\mathbf{v}_{\mathrm{b}})\rvert$ (solid red) reaches 0.20–0.34 in early training. The ${\sim}40\times$ gap demonstrates that the backbone emerges from optimizer integration, not from instantaneous gradient structure. The signed update alignment (dashed red) reveals a persistent negative bias that flips sign near step 5000.
Sign structure.
The signed cosine $C(t)$ is persistently negative throughout early training (the optimizer drifts in the $\mathbf{v}_{\mathrm{b}}$ direction), then flips to positive around step 5000–5200. Defining the sign-flip step $t^{*}$ as the first checkpoint where $C(t)$ crosses zero:
Table 4: Sign flip in update–backbone alignment.
Seed $t^{*}$ Early $C(t)$ (Blk 0) Late $C(t)$ (Blk 0)
42 ${\sim}5200$ 0.27-0.27 +0.11+0.11
271 ${\sim}5000$ 0.29-0.29 +0.07+0.07
The sign reversal coincides with the $\lambda$-transition (step 4000, $\lambda:2\to 4$) and the onset of oscillation damping. The reversal is present in both seeds (Table˜4).
4.4 Block Localization: The First Transformer Block Drives the Backbone
Which blocks contribute the gradient bias that creates the backbone? At each checkpoint, we compute the signed projection of the combined-loss gradient onto the backbone, per block $\ell$:
$b_{\ell}(t)\;=\;\langle\mathbf{g}_{\ell}(t),\,\mathbf{v}_{\mathrm{b}}^{(\ell)}\rangle,$ (24)
where $\mathbf{g}_{\ell}(t)$ is the gradient restricted to block $\ell$ and $\mathbf{v}_{\mathrm{b}}^{(\ell)}$ is the block-$\ell$ backbone.

### v2

# arXiv:2602.23696v2[cs.LG] 02 Mar 2026

@@ Abstract
We analyze cumulative parameter trajectories of transformer training under AdamW and identify a dominant low-dimensional drift direction (“backbone”) that captures 60–80% of long-horizon displacement from initialization. This direction is highly stable over rolling training windows yet reorients gradually across phases, particularly following objective reweighting. Per-batch gradients exhibit near-noise-floor alignment with the backbone, whereas optimizer-integrated updates align strongly with it, indicating that the structure emerges from accumulated optimizer dynamics rather than instantaneous gradient geometry.
Replacing AdamW with SGD-family optimizers eliminates this structure, and reducing $\beta_{2}$ smoothly degrades backbone dominance and reheating recoverability. Reheating experiments show that transverse probe modes can be transiently re-excited without substantially altering accumulated backbone drift.
These results provide a trajectory-level characterization of optimizer-induced geometric structure in transformer training and shift attention from instantaneous gradient properties to cumulative update dynamics.
@@ 1 Introduction
In this work, we study the global parameter trajectory of transformer training under AdamW (Loshchilov and Hutter, 2019). Rather than analyzing local curvature or single-step gradients, we examine cumulative displacement from initialization across checkpoints. We find that training admits a dominant low-dimensional drift direction—which we term the backbone—that captures the majority of cumulative parameter movement. Across blocks and seeds, the first principal component of uncentered trajectory PCA explains 60–80% of total drift. This direction is highly stable over rolling training windows yet reorients gradually across longer horizons, particularly following objective reweighting.
Crucially, the backbone is not aligned with instantaneous gradient directions and is nearly orthogonal to leading Fisher curvature modes (Martens, 2020). Per-batch gradients are close to isotropic relative to the backbone. However, the optimizer-integrated update—after momentum accumulation and adaptive per-parameter normalization (Kingma and Ba, 2015)—exhibits strong alignment with it. This indicates that the backbone is not a property of the loss landscape alone, but emerges from accumulated optimizer dynamics.
@@ 3.2 Temporal Stability of the Backbone
The rolling-window backbone direction is highly stable locally: across adjacent windows, $c(w)\approx 0.997$$0.998$ for both seeds, indicating extremely small curvature over 2000-step horizons. However, when comparing phase-level backbones estimated over early (0–4k) and late (4k–10k) intervals, $\lvert\langle v_{E},\,v_{L}\rangle\rvert\approx 0.32$, revealing substantial long-horizon reorientation ($\approx 71^{\circ}$). Thus the backbone is locally stable but globally curved—it is not a fixed direction but a smooth, slowly rotating tangent to a curved one-dimensional manifold in parameter space.
@@ 3.3 Backbone–Residual Decomposition
4 Mechanism: Optimizer-Induced Slow Manifold and Controlled Rotation
4.1 Global Backbone Direction and Spectral Concentration
We begin with the global structure of the cumulative parameter trajectory. For each seed, we flatten all trunk parameters (attention QKV/output projections and MLP weights; D=25,165,824D=25{,}165{,}824 parameters) at each checkpoint $t\in\{200,400,\ldots,10000\}$, compute drifts $\Delta\theta_{t}=\theta_{t}-\theta_{1}$, row-normalize, and extract the top right singular vector $v_{b}$ via SVD.
Across the full 10k trajectory, PC1 captures a dominant fraction of row-normalized variance:
Seed 42 Seed 271
PC1 (row-normalized) 71.4% 71.7%
PC2 (row-normalized) 12.1% 11.8%
PC1 (raw, unnormalized) 79.1% 79.3%
Trunk dimension $D$ 25,165,824
The trajectory is dominated by a single direction, but as we show below, this direction is not static.
4.2 Local Tangent Stability
Let $v_{\mathrm{roll}}(t)$ denote the first principal component of the row-normalized, uncentered drift matrix computed in a rolling window of $W=10$ checkpoints ($\approx$2000 steps) centered at step $t$. To measure local curvature, define adjacent-window alignment:
$\rho(t)=\bigl|\langle v_{\mathrm{roll}}(t),\;v_{\mathrm{roll}}(t+\Delta)\rangle\bigr|.$
Across the full training run (both seeds):
Seed 42 Seed 271
$\mathbb{E}[\rho(t)]$ 0.800 0.793
$\max\;\rho(t)$ 0.936 0.938
$\min\;\rho(t)$ 0.749 0.702
Step at $\min\;\rho$ $\approx$5000 $\approx$5000
Adjacent windows maintain high alignment ($\rho>0.7$ everywhere), confirming that the backbone tangent varies smoothly. However, $\rho$ is not uniformly close to 1: the dip near step 5000 signals a region of concentrated curvature.
4.3 Cumulative Turning and Global Reorientation
Local stability does not imply global rigidity. Small rotations accumulate. To quantify long-horizon reorientation, we compute phase-level backbones via global SVD over disjoint intervals:
  • $v_{E}$: PC1 over the early interval (steps 1–4000, relative to $\theta_{1}$),
  • $v_{L}$: PC1 over the late interval (steps 4000–10000, relative to $\theta_{4000}$).
These phase backbones are substantially misaligned:
Seed 42 Seed 271
PC1 (early, 0–4k) 60.0% 59.2%
PC1 (late, 4k–10k) 81.3% 81.7%
$|\langle v_{E},v_{L}\rangle|$ 0.323 0.323
The cosine similarity of 0.32 corresponds to an angle of $\approx 71^{\circ}$. Thus:
  • Locally: the backbone tangent is smooth.
  • Globally: the slow direction rotates by $\approx 71^{\circ}$ over the full run.
The cumulative trajectory lies on a smooth but substantially curved one-dimensional manifold in parameter space.
Notably, the late phase has stronger rank-1 concentration (PC1 $\approx 81\%$) than the early phase (PC1 $\approx 60\%$). After the $\lambda$-switch, the optimizer locks into a tighter low-rank drift — but in a different direction.
4.4 Transition Region and $\lambda$-Induced Bending
To track how reorientation unfolds, we measure the alignment of each sliding-window backbone to the two phase backbones:
$A_{E}(t)=\bigl|\langle v_{\mathrm{roll}}(t),\;v_{E}\rangle\bigr|,\qquad A_{L}(t)=\bigl|\langle v_{\mathrm{roll}}(t),\;v_{L}\rangle\bigr|.$
These reveal a clean handoff (representative values from seed 42; seed 271 is nearly identical):
Window center $A_{E}$ $A_{L}$ Phase
1100 0.83 0.16 Early backbone dominant
2300 0.61 0.28 Fading
3500 0.02 0.16 Dead zone — neither backbone
4700 0.21 0.59 Late backbone emerging
5300 0.20 0.69 Peak late alignment
7100 0.18 0.29 Late backbone fading
8900 0.18 0.02 Neither — late plateau
Around step 3500, $A_{E}\approx 0$ and $A_{L}\approx 0.15$: the optimization trajectory passes through a geometric transition zone aligned to neither phase backbone. Late-phase alignment $A_{L}$ peaks near step 4700–5300, then decays. By step 8000+, the drift direction is orthogonal to both $v_{E}$ and $v_{L}$, consistent with the very late plateau phase.
The rotation dip in $\rho(t)$ (minimum at step $\approx$5000) coincides with this transition, confirming that the $\lambda$-switch (step 4000) perturbs the training vector field and bends the slow manifold. Importantly, the bending is continuous: there is no discontinuity in tangent direction.
4.5 Power-Law Dynamics: Before and After the $\lambda$-Switch
The backbone coordinate $a(t)=\langle\Delta\theta_{t},v_{b}\rangle$ and residual $\|r(t)\|=\|\Delta\theta_{t}-a(t)v_{b}\|$ obey distinct power-law regimes. We fit $|a(t)|=C_{a}t^{\gamma_{a}}$ and $\|r(t)\|=C_{r}t^{\gamma_{r}}$ in four windows:
$\gamma_{a}$ (backbone) $\gamma_{r}$ (residual)
Regime Seed 42 Seed 271 Seed 42 Seed 271
0–2000 +2.07+2.07 +2.07+2.07 +1.24+1.24 +1.26+1.26
2000–4000 +0.67+0.67 +0.67+0.67 0.39-0.39 0.43-0.43
4000–6000 +0.15+0.15 +0.12+0.12 1.33-1.33 1.38-1.38
6000–10000 0.19-0.19 0.20-0.20 +0.36+0.36 +0.32+0.32
All fits have $R^{2}>0.85$ (most >0.97>0.97). Three dynamical phases emerge:
  1. 1.
    Acceleration (0–2k): Backbone grows as $a\sim t^{2.1}$, residual as $r\sim t^{1.2}$. Both parallel and perpendicular motion accelerate.
  2. 2.
    Consolidation (2k–4k): Backbone decelerates ($\gamma_{a}\approx 0.67$) while residual actively contracts ($\gamma_{r}\approx-0.4$). The model consolidates along the backbone.
  3. 3.
    Post-switch: The $\lambda$-switch at step 4000 triggers the sharpest residual collapse ($\gamma_{r}\approx-1.35$ in 4k–6k), followed by backbone retreat ($\gamma_{a}\approx-0.19$) and residual re-growth ($\gamma_{r}\approx+0.34$) in the 6k–10k plateau.
The overall before/after picture: $\gamma_{a}$ drops from +1.74+1.74 (0–4k) to 0.08-0.08 (4k–10k), and $\gamma_{r}$ drops from +0.84+0.84 to 0.31-0.31, confirming that the $\lambda$-switch arrests and reverses the dominant drift.
4.6 Correlation Between Probe Accuracy and Residual Geometry
To connect backbone geometry to task performance, we compute Pearson correlations between probe OOD accuracy $p_{\mathrm{ood}}$ and residual norm $\|r(t)\|$:
Regime Seed 42 Seed 271
Full range (0–10k) +0.61+0.61 +0.45+0.45
0–4000 +0.85+0.85 +0.76+0.76
4000–10000 +0.43+0.43 +0.28+0.28
6000–10000 0.73-0.73 0.79-0.79
The correlation flips sign: in the early phase, growing residual energy accompanies improving probe accuracy ($r\approx+0.8$). After step 6000, the re-growing residual is associated with declining probe accuracy ($r\approx-0.76$). The early residual reflects productive exploration of directions useful for the probe task; the late residual reflects drift away from the solution.
4.7 Optimizer Integration Creates the Slow Direction
The effective update under AdamW is
$u_{t}=-\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}-\mu\,\theta_{t},$
where momentum integrates gradient history and second-moment normalization rescales coordinates. Two mechanisms produce coherent slow drift:
(i) Momentum integrates weak signed bias.
Even if instantaneous projections onto $v_{b}(t)$ are small, a persistent sign bias accumulates in the momentum buffer. Momentum acts as a temporal low-pass filter, increasing the signal-to-noise ratio of temporally coherent gradient components.
(ii) Adaptive normalization suppresses incoherent variance.
Coordinates with high variance but low mean are downscaled; coordinates with small but consistent bias are preserved. This selectively amplifies temporally coherent directions.
Evidence from the $\beta_{2}$ ablation supports this mechanism. Disabling the second-moment estimator ($\beta_{2}=0$) dramatically weakens backbone concentration and update alignment:
$\beta_{2}$ PC1% Drift magnitude Mean $|\cos(u,v_{b})|$ Best $p_{\mathrm{ood}}$
0.99 68.1 106 0.226 0.951
0.95 68.4 108 0.225 0.939
0.90 66.3 113 0.220 0.814
0.80 63.4 128 0.214 0.682
0.0 51.6 211,694 0.099 0.005
@@ 4.7 Optimizer Integration Creates the Slow Direction
4.8 Update–Backbone Alignment and Sign Reversal
Using 200-step cumulative updates $u(t)=\theta(t)-\theta(t-200)$, we observe strong alignment with the global backbone (Figure˜2). For seed 271 (layer 0, representative):
$|\cos\angle(u(t),v_{b})|\approx 0.15\text{--}0.32\quad\text{(steps 100--1900)},$
@@ 4.8 Update–Backbone Alignment and Sign Reversal
The signed alignment exhibits a characteristic reversal:
  • Steps 100–4700: $\cos<0$ (drift along $v_{b}$), peaking at $|\cos|\approx 0.32$ near step 1700.
  • Step $\approx$5100: sign flip ($\cos$ crosses zero).
  • Steps 5100–10000: $\cos>0$ (drift along $v_{b}$), saturating at $|\cos|\approx 0.11$.
This sign reversal coincides with both the $\lambda$-switch and the backbone rotation documented in §4.4. The optimizer does not merely slow; it redirects cumulative drift. The post-reversal alignment is weaker ($0.11$ vs. $0.32$), consistent with the slower late-phase dynamics.
Figure 2: Gradient vs. optimizer-update alignment with the backbone (Block 0, seed 42). The 200-step optimizer update $\mathbf{u}_{t}$ (blue) aligns strongly with the backbone ($|\cos|\approx 0.15$$0.34$), peaking before the $\lambda$-transition (dashed line) and declining afterward. Per-batch gradients $\mathbf{g}_{t}$ (grey squares) remain at the random noise floor (${\sim}4\times 10^{-4}$) throughout. The backbone emerges from optimizer integration, not instantaneous gradient structure.
4.9 Backbone Stiffening Under Rotation
Fisher information analysis shows that curvature along the backbone direction increases by orders of magnitude during training (Figure˜3). The Rayleigh quotient $q_{b}=v_{b}^{T}Hv_{b}$ (approximated via Fisher diagonal) and anisotropy ratio $\alpha=q_{b}/\mathbb{E}[q_{\mathrm{rand}}]$ evolve as follows (seed 42):
Step Label $q_{b}$ Anisotropy $\alpha$
200 Init $2.4\times 10^{-6}$ 1.35
1800 Peak (pre-switch) $2.5\times 10^{-6}$ 1.92
4800 Trough (post-switch) $1.6\times 10^{-4}$ 12.4
9600 Late peak $8.1\times 10^{-3}$ 4.81